# Modeling the TMD shape function in $J/\psi$ electroproduction

### v2

@@ 5 Conclusions
While this work was in preparation, a related study appeared in Ref. Maxia:2025zee . It is important to stress that the two works differ in how the shape function has been described, as we explained in the main text. Phenomenological predictions in Ref. Maxia:2025zee follow the matching between fixed order and TMD calculations, providing a crucial step toward a complete picture of the cross-section. In contrast, and unlike this study, in our analysis we explicitly implement the calculation of the hard part at NLO. Furthermore, our work focuses on the phenomenological implementation of the TMDShF described at the operator level in resummed cross-section predictions. Our NNLL resummed predictions, supplemented by a consistent non-perturbative parametrization, yield narrower uncertainty bands and enhanced stability compared to NLL, underscoring the necessity of precision resummation and NP modeling in quarkonium TMD phenomenology.
@@ Acknowledgements.
We are grateful to Melih A. Ozcelik and Shubhendu P. Mandal for helpful discussions. This work was supported by Ikerbasque (Basque Foundation for Science), by the Basque Government through the grant IT1628-22, and by the Spanish State Research Agency through the grants PID2022-136510NB-C33 and CNS2022-135186.
@@ Appendix A Notation and formulae
The Born cross-sections for $J/\psi$ photo- and electroproduction can be obtained with the so-called projectors method (see, e.g., Ref. Maltoni:1997pt ) or by using the matching tensors (see Eq. (6)). To present the results in a clean way, we write the Born cross-section as
$\sigma_{[n]\text{Born}}=\frac{4\pi^{2}\alpha_{em}^{2}\alpha_{s}e_{Q}^{2}}{y\,Q^{2}M(M^{2}+Q^{2})}\;\frac{\hat{\sigma}_{[n]}}{N_{pol}}$ (37)
where the factors $\hat{\sigma}_{[n]}$ are gathered in the Table 3. The parameter $N_{pol.}$ refers to the number of polarizations in $d=4-2\epsilon$ dimensions: $N_{J=0}=1$, $N_{1}=(3-2\epsilon)$ and $N_{2}=(5-2\epsilon)(1-\epsilon)$. The parameter $N_{col.}$ is the color normalization factor. It was proven that these results coincide with those presented in Ref. Bacchetta:2018ivt .
@@
As mentioned in the text, the finite part of the TMD shape function anomalous dimension was obtained in Ref. Echevarria:2024idp :
@@
$\begin{gathered}\gamma_{H}=2\,\Gamma_{cusp}(\alpha_{s})\,\hbox{ln}\left(\frac{f(M^{2},Q^{2})}{\mu_{H}^{2}}\right)+2\,\gamma_{nc}(\alpha_{s})\;,\\ \text{with}\qquad\gamma_{nc}=-\frac{\alpha_{s}C_{A}}{2\pi}\;.\end{gathered}$ (40)
@@
After solving the renormalization-group-evolution equations for the LDMEs, one gets the following expression, e.g., see the Ref. Butenschoen:2020mzi :
@@
As we mentioned in Eq. (11), the TMDShF can be written in terms of their collinear analogous through the following matching coefficients, derived in Ref. Echevarria:2024idp :
@@ Appendix B One-loop virtual corrections
As mentioned in the main text, in order to calculate the partonic cross-section for the process $\gamma^{+}g\to c\bar{c}$ and thereby extract the NLO hard contribution, we have followed the procedure used for the case of a real photon, as presented in the Ref. Maltoni:1997pt . In this appendix, we summarize the most relevant aspects of the computation and present partial results, aiming to provide the greatest possible level of detail.
@@
$\begin{gathered}q^{2}=-Q^{2},\quad p_{g}^{2}=0,\quad p^{2}=\bar{p}^{2}=m^{2},\quad P_{\psi}^{2}=4m^{2},\\ q\cdot p=q\cdot\bar{p}=\frac{4m^{2}-Q^{2}}{4}\quad\text{and}\quad p_{g}\cdot p=p_{g}\cdot\bar{p}=\frac{4m^{2}+Q^{2}}{4}\;.\end{gathered}$ (45)
@@
We would like to clarify how we handle the Coulomb singularity in our calculation, because unlike the result for photoproduction Maltoni:1997pt , we do not explicitly obtain the Coulomb pole $1/\text{v}$ in Eq (7).
@@
Note we do not obtain any dependence in the relative velocity v, as is obvious when setting $q=0$. However, the result reported in Ref. Maltoni:1997pt for this scalar integral contains an additional term inside the brackets of the previous expression, namely the Coulomb pole $\pi^{2}/2\text{v}$. This arises from considering $p\neq\bar{p}$ with $p\cdot\bar{p}=m^{2}\bigl(\tfrac{1+\text{v}^{2}}{1-\text{v}^{2}}\bigr)$
@@
    $-\frac{\text{CF}\left(64\rho^{6}+100\rho^{5}+116\rho^{4}+221\rho^{3}+546\rho^{2}+421\rho+96\right)}{2(\rho+1)(2\rho+1)^{2}\left(\rho^{2}-3\rho+6\right)}\;,$
@@
    $f_{6}$ $=\frac{2C_{F}\left(-2\rho^{3}+4\rho^{2}+7\rho+3\right)-C_{S}\left(2\rho^{3}-4\rho^{2}+19\rho+27\right)}{2(\rho+1)\left(\rho^{2}-3\rho+6\right)}\;,$

### v3

@@ 5 Conclusions
While this work was in preparation, a related study appeared in Ref. Maxia et al. (2025). It is important to stress that the two works differ in how the shape function has been described, as we explained in the main text. Phenomenological predictions in Ref. Maxia et al. (2025) follow the matching between fixed order and TMD calculations, providing a crucial step toward a complete picture of the cross-section. In contrast, and unlike this study, in our analysis we explicitly implement the calculation of the hard part at NLO. Furthermore, our work focuses on the phenomenological implementation of the TMDShF described at the operator level in resummed cross-section predictions. Our NNLL resummed predictions, supplemented by a consistent non-perturbative parametrization, yield narrower uncertainty bands and enhanced stability compared to NLL, underscoring the necessity of precision resummation and NP modeling in quarkonium TMD phenomenology.
@@ Acknowledgements.
We are grateful to Melih A. Ozcelik and Shubhendu P. Mandal for helpful discussions. This work was supported by Ikerbasque (Basque Foundation for Science), by the Basque Government through the grant IT1628-22, and by the Spanish State Research Agency through the grants PCI2022-132984, PID2022-136510NB-C33 and CNS2022-135186.
@@ Appendix A Notation and formulae
The Born cross-sections for $J/\psi$ photo- and electroproduction can be obtained with the so-called projectors method (see, e.g., Ref. Maltoni et al. (1998)) or by using the matching tensors (see Eq. (6)). To present the results in a clean way, we write the Born cross-section as
$\sigma_{[n]\text{Born}}=\frac{4\pi^{2}\alpha_{em}^{2}\alpha_{s}e_{Q}^{2}}{y\,Q^{2}M(M^{2}+Q^{2})}\;\frac{\hat{\sigma}_{[n]}}{N_{pol}N_{col}}$ (38)
where the factors $\hat{\sigma}_{[n]}$ are gathered in the Table 3. The parameter $N_{pol.}$ refers to the number of polarizations in $d=4-2\epsilon$ dimensions: $N_{J=0}=1$, $N_{1}=(3-2\epsilon)$ and $N_{2}=(5-2\epsilon)(1-\epsilon)$. The parameter $N_{col}$ is the color normalization factor. It was proven that these results coincide with those presented in Ref. Bacchetta et al. (2020).
@@
As mentioned in the text, the finite part of the TMD shape function anomalous dimension was obtained in Ref. Echevarria et al. (2024):
@@
$\begin{gathered}\gamma_{H}=2\,\Gamma_{cusp}(\alpha_{s})\,\hbox{ln}\left(\frac{f(M^{2},Q^{2})}{\mu_{H}^{2}}\right)+2\,\gamma_{nc}(\alpha_{s})\;,\\ \text{with}\hskip 16.38895pt\gamma_{nc}=-\frac{\alpha_{s}C_{A}}{2\pi}\;.\end{gathered}$ (41)
@@
After solving the renormalization-group-evolution equations for the LDMEs, one gets the following expression, e.g., see the Ref. Butenschoen and Kniehl (2020):
@@
As we mentioned in Eq. (11), the TMDShF can be written in terms of their collinear analogous through the following matching coefficients, derived in Ref. Echevarria et al. (2024):
@@ Appendix B One-loop virtual corrections
As mentioned in the main text, in order to calculate the partonic cross-section for the process $\gamma^{+}g\to c\bar{c}$ and thereby extract the NLO hard contribution, we have followed the procedure used for the case of a real photon, as presented in the Ref. Maltoni et al. (1998). In this appendix, we summarize the most relevant aspects of the computation and present partial results, aiming to provide the greatest possible level of detail.
@@
$\begin{gathered}q^{2}=-Q^{2},\qquad p_{g}^{2}=0,\qquad p^{2}=\bar{p}^{2}=m^{2},\qquad P_{\psi}^{2}=4m^{2},\\ q\cdot p=q\cdot\bar{p}=\frac{4m^{2}-Q^{2}}{4}\qquad\text{and}\qquad p_{g}\cdot p=p_{g}\cdot\bar{p}=\frac{4m^{2}+Q^{2}}{4}\;.\end{gathered}$ (46)
@@
We would like to clarify how we handle the Coulomb singularity in our calculation, because unlike the result for photoproduction Maltoni et al. (1998), we do not explicitly obtain the Coulomb pole $1/\text{v}$ in Eq (7).
@@
Note we do not obtain any dependence in the relative velocity v, as is obvious when setting $q=0$. However, the result reported in Ref. Maltoni et al. (1998) for this scalar integral contains an additional term inside the brackets of the previous expression, namely the Coulomb pole $\pi^{2}/2\text{v}$. This arises from considering $p\neq\bar{p}$ with $p\cdot\bar{p}=m^{2}\bigl(\tfrac{1+\text{v}^{2}}{1-\text{v}^{2}}\bigr)$
@@
    $-\frac{C_{F}\left(64\rho^{6}+100\rho^{5}+116\rho^{4}+221\rho^{3}+546\rho^{2}+421\rho+96\right)}{2(\rho+1)(2\rho+1)^{2}\left(\rho^{2}-3\rho+6\right)}\;,$
@@
    $f_{6}$ $=\frac{2C_{F}\left(-2\rho^{3}+4\rho^{2}+7\rho+3\right)-C_{A}\left(2\rho^{3}-4\rho^{2}+19\rho+27\right)}{2(\rho+1)\left(\rho^{2}-3\rho+6\right)}\;,$